# TLDP: An Algorithm of Local Differential Privacy for Tensors

### v1

@@ Abstract
Tensor-valued data, increasingly common in applications like spatiotemporal modeling and social networks, pose unique challenges for privacy protection due to their multidimensional structure and the risk of losing critical structural information. Traditional local differential privacy (LDP) methods, designed for scalars and matrices, are insufficient for tensors, as they fail to preserve essential relationships among tensor elements. We introduce TLDP, a novel LDP algorithm for Tensors, which employs a randomized response mechanism to perturb tensor components while maintaining structural integrity. To strike a better balance between utility and privacy, we incorporate a weight matrix that selectively protects sensitive regions. Both theoretical analysis and empirical findings from real-world datasets show that TLDP achieves superior utility while preserving privacy, making it a robust solution for high-dimensional data.
@@
Article submission, IEEE, IEEEtran, journal, , paper, template, typesetting.
@@ I Introduction
Differential privacy is an algorithm that anonymizes original data while processing large datasets, thereby obscuring sensitive information and reducing the risk of data inference or re-identification. In recent years, it has gained widespread adoption in tasks such as frequency estimation [1] and count estimation [2]. This privacy protection mechanism typically relies on trusted third-party servers, which can present implementation challenges in real-world scenarios. As concerns over the protection of sensitive information grow, such as location data and social media communications, Local Differential Privacy (LDP) is becoming increasingly prevalent. LDP decentralizes privacy protection by enabling users to directly handle their data, allowing parties to collect aggregate information while safeguarding individual privacy without relying on a trusted intermediary. There are various methods for implementing LDP for scalar raw data, the most common of which is the randomized response mechanism [3]. This technique has been widely used since its inception. For example, researchers at Google developed a tool known as RAPPOR [4, 5], which uses a randomized response method and has been integrated into Chrome. This innovative tool enables Google to collect data in response to various queries about the browser’s settings, including the default search engine and homepage, with the goal of identifying and understanding instances of unwanted or malicious changes to user configurations. To improve accuracy and reduce communication costs, other enhanced LDP protocols have been proposed successively, including Direct Encoding (DE) [6], Unary Encoding (UE) [4], Histogram Encoding (HE) [4], and Local Hashing (LH) [7, 8].
However, in many applications, such as spatiotemporal social network analysis, user behavior modeling, and large-scale data processing, data is often stored in the form of tensors. Tensors, which generalize matrices to an arbitrary number of dimensions, are well-suited to represent diverse data types, including multimedia data, graph data, and data cubes. Treating tensors merely as collections of scalars risks losing critical structural information and interrelationships among elements. For instance, consider a symmetric tensor: if each element is perturbed using traditional methods, such as adding independent and identically distributed noise, the probability that the tensor still remains symmetric is extremely low. Similarly, flattening a data cube into vectors would result in the loss of significant information. Moreover, tensors present greater challenges than scalars, even with the same privacy budget $(\epsilon)$, due to their multidimensional nature. The added complexity of extra dimensions complicates the privacy considerations, making them more intricate than those for scalar data. This complexity is further compounded by the fact that tensors with different shapes can yield different data interpretations, highlighting the need for a more sophisticated approach to privacy protection. A thorough analysis is essential to ensure that privacy concerns are adequately addressed, especially given the potential for varied and complex data representations within tensor frameworks.
Designing LDP mechanisms for tensor-valued data presents unique difficulties, arising not only from the complex nature of high-dimensional distributions but also from the lack of universal privacy frameworks that accommodate diverse data types. While LDP mechanisms have been developed for scalars [9], vectors, and matrices [10], a unified mechanism capable of handling various data forms remains unavailable. In this work, we aim to address this gap by proposing a robust, unified framework for LDP that is adaptable to the specific structure of tensors.
We introduce TLDP, a Local Differential Privacy algorithm designed for Tensors of varying dimensions, expanding the scope of LDP applications. Several approaches have recently been proposed for tensor data, such as adding noise to the query function to obtain differential privacy [11], and applying randomized responses to individual inputs [12]. However, the former method requires adding more noise under LDP than the traditional Differential Privacy (DP) technique, which can degrade data utility due to excessive noise. The latter approach, while protecting individual inputs in the process of Candecomp/Parafac factorization by perturbing factor matrices through randomized response, still carries the risk of leaking original data with some probability, posing a privacy concern. In light of these issues, we propose an innovative approach that adds noise to tensor data by randomly selecting components of the tensor with a randomized response mechanism to accomplish local differential privacy. Moreover, recognizing that different parts of a tensor may have varying sensitivities, we incorporate a weight matrix into the algorithm. For instance, when processing image data as sensitive information, assigning higher weights to more sensitive regions can improve data utility while minimizing the privacy budget.
@@
  • We propose an algorithm that ensures the local differential privacy of tensor-valued data through randomized response. This approach not only preserves users’ sensitive information but also enables the perturbed data to be effectively used for training models.
  • We further enhance the algorithm by incorporating a weight matrix, which selectively reduces the impact of perturbation on important portions of the data, thereby strengthening privacy protection.
  • We provide theoretical proof of privacy under the LDP mechanism and present a series of experiments performed on various tensor-valued models and datasets. The results demonstrate that our method achieves superior utility to other mechanisms while maintaining the same level of privacy.
@@ II Related Work
II-A Randomized Response
In sample surveys, different questions may elicit different degrees of cooperation from respondents. People are generally more willing to share their genuine opinions on innocent topics, such as “What is your favorite book/movie?”, but are more likely to avoid or provide misleading answers when confronted with revealing questions, such as “Do you have a criminal record?”. To mitigate this evasive response bias, Stanley L. Warner [3] introduced the randomized response, a survey technique specifically designed to preserve respondent privacy, in 1965, commonly referred to as $w$-RR. In this method, respondents are randomly assigned to one of two groups with probability $p$ and $1-p$ and are then instructed to answer “yes” or “no” to a question based on their group assignment. The key innovation of this technique is its ability to protect individual privacy while still enabling the collection of meaningful data. Since interviewers only receive the aggregate count of ”yes” and ”no” responses, they cannot ascertain the specific question to which each answer corresponds, thus maintaining anonymity and promoting truthful answers. This approach has been widely adopted in surveys involving sensitive topics, ensuring the confidentiality of respondents while providing valuable insights into the population’s attitudes and behaviors.
The $w$-RR technique is specifically designed for discrete datasets with exactly two distinct values. To extend its applicability, two primary paths for improvement can be explored. The first involves encoding and transforming the variable’s multiple values, as demonstrated in methods such as Randomized Aggregatable Privacy-Preserving Ordinal Response (RAPPOR) [4] and S-Hist [7], to ensure compatibility with the binary nature expected by $w$-RR.
The RAPPOR employs Bloom filter and hash functions to encode a single element into a vector, then perturb the origin value through both permanent and instantaneous randomized responses. It represents a method for collecting statistical data from end-user applications while preserving anonymity and ensuring robust privacy protections.
@@ II-A Randomized Response
However, RAPPOR has two main drawbacks: (1) the transmission overhead between the user and the data aggregator is relatively high, as each user is required to send a vector whose length is determined by the size of its bloom filter; (2) the data collector must pre-collect a list of candidate strings for frequency counting. To address the first issue, the S-Hist method [7] offers a solution in which each user encodes the string, randomly selects a bit, applies random response techniques to perturb it, and then sends it to the data collector. This approach significantly reduces the transmission cost by minimizing the amount of data each user needs to send. To overcome the second limitation, one approach is to improve the distribution of the $w$-RR technique, making it applicable to variables with more than two values. This is achieved through mechanisms such as $k$-RR [13] and $O$-RR [14]. In the $k$-RR method, the original data is preserved with probability $p$ and is flipped to any other value with a probability of $\frac{1-p}{n-1}$, where $n$ is the number of possible values. This mechanism ensures that the perturbation is uniformly distributed across all possible values, thereby preserving the privacy of the data. Building on this principle, Kairouz et al. [14] further proposed the $O$-RR method, which handles cases where the possible values of a variable are unknown by introducing hash mapping and grouping operations. Hash mapping allows the method to focus on the encoded values rather than the original strings, eliminating the need to pre-collect a list of candidate strings. Additionally, the use of grouping operations reduces the probability of hash mapping value collisions. This combination of techniques enhances the robustness and flexibility of differential privacy in various data collection scenarios.
@@
II-B Noise Mechanisms
In conventional differential privacy approaches, data perturbation is achieved through various noise mechanisms, some of which are independent of each other. These include the Gaussian mechanism [18], Laplace mechanism [19], and Matrix Variate Gaussian (MVG) [20].
Although the mechanisms mentioned above do not directly apply to tensor-valued queries, our research is nevertheless grounded in the fundamental principle of preserving privacy through the injection of Gaussian or Laplace noise. The Gaussian mechanism introduces independent and identically distributed (i.i.d.) Gaussian noise, calibrated according to the $l_{2}$-sensitivity of the query, guaranteeing $(\epsilon,\delta)$-differential privacy. Similarly, By incorporating noise derived from the Laplace distribution, scaled according to the query function’s $l_{1}$-sensitivity, the Laplace mechanism enforces strong $\epsilon$-differential privacy. Building upon these concepts, our mechanism is specifically tailored to address the complexities of tensor data. We demonstrate that our mechanism attains $\epsilon$-differential privacy for tensors, as detailed in Section III. This innovation bridges the gap between traditional differential privacy mechanisms and the sophisticated requirements of high-dimensional data, preserving privacy while maintaining utility. The Multivariate Gaussian (MVG) mechanism is tailored for matrix-valued queries and incorporates matrix-valued noise to ensure $(\epsilon,\delta)$-differential privacy. The $l_{2}$-sensitivity of the MVG mechanism is measured using the Frobenius norm, which quantifies the discrepancy between two neighboring matrices.
Mechanisms that derive their privacy guarantees from the above mechanisms are well-established, as exemplified by the composition schemes referenced in [18, 21, 22, 23, 24, 25], as well as privacy amplification through sampling as outlined in [26]. Among these, [21] presents an innovative accounting approach for the Gaussian mechanism that minimizes the total additive noise while preserving the same level of privacy protection. In contrast, Kairouz et al. [13] propose an optimal composition scheme that is applicable to a broader range of noise distributions. The studies in [23, 25] provide dynamic accounting methods that adjust based on the algorithm’s runtime convergence. Furthermore, [24] ensures a global differential privacy guarantee in environments where datasets are continuously expanding.
II-C Learning With Differential Privacy
A range of studies have explored the integration of differential privacy techniques into machine learning models, addressing various privacy-preserving objectives. These efforts encompass differentially private data inputs [20, 27], outputs [28, 29], gradients [23, 24, 25, 30, 31, 32, 33], and even objective functions [34, 35, 36]. Our focus is especially on machine learning applications, which often involve high-dimensional or complex data. The proposed mechanism is both robust and adaptable, designed to protect inputs, outputs, and gradients, especially when the data is structured in tensor format, as evidenced by the result of the final simulation experiments.
II-D Summaries
While research on noise mechanisms and random response systems has advanced significantly in isolation over the years, the potential for integrating the two approaches has remained largely underexplored. To mitigate the effect of tensor dimensions on noise levels, we introduce a novel method that applies the randomized response mechanism to select noise positions rather than perturbing the data directly. This inherent randomness substitutes for traditional noise, effectively thwarting attackers from inferring the original data.
@@ III Preliminaries
In this section, we aim to provide our readers with foundational knowledge to facilitate a smoother comprehension of our work.
@@ III-A Local Differential Privacy
These years people have become increasingly cautious about their privacy and often feel compelled to share their personal information to benefit from customized services, including personalized searches and recommendations. While the sharing of data offers mutual benefit to both data providers and data analysts, individuals also desire plausible deniability concerning their sensitive information, so it is imperative to protect users’ privacy against centralized servers. LDP, an extension of centralized differential privacy techniques, introduces key features that provide two primary advantages: (1) Quantification of Privacy Protection: these techniques account for the background knowledge that an adversary might possess, providing a quantifiable measure of privacy protection. This ensures that privacy guarantees are not merely theoretical but can be precisely calculated and adjusted based on the level of potential threat. (2) Local Data Perturbation: by introducing local perturbations to the data at the point of collection or usage, rather than relying on a centralized system, this technique enhances the privacy of individuals against potential privacy invasions by untrusted third parties. This aligns with the growing need for privacy in the age of the Internet. In this paper, we assume that each user has a tensor value $\mathcal{X}$, representing their sensitive information.
Definition 1 (LDP).
An algorithm M satisfies $\epsilon$-LDP ($\epsilon$-LDP), where $\epsilon\geq 0$ , if and only if for any tensor-value input $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, we have
$\forall y\in Range(\textbf{M}),Pr[\textbf{M}(\mathcal{X}_{1})=y]\leq e^{\epsilon}Pr[\textbf{M}(\mathcal{X}_{2})=y],$
According to this definition, a single bit perturbed by the randomized response mechanism [3] satisfies $\ln(\frac{p}{1-p})$-LDP.
Problem Definition and Notations. We suppose that in this article, each user has a tensor-value $\mathcal{X}\in\mathbf{R}^{I_{1}\times I_{2}\times\dots\times I_{N}}$. We utilize $\Delta$ to represent the size of the domain for the values of each component of the tensor and $[\Delta]$ to represent the set $\{1,2,\dots,\Delta\}$. Without compromising generality, we assume the input domain is $[\Delta]$, $i.e.\mathcal{X}_{i_{1}i_{2}\dots i_{N}}\in[\Delta]$. If the tensor takes continuous values, it is recommended to discretize it initially.
The notations used in this paper are summarized in Table I for easy reference.
@@
$I$ Product of the tensor’s dimensions
@@
$\Delta$ Range of the initial data
@@
$W$ Weighted matrix
Our goal is to perturb the original tensor before transmitting it to the untrusted central server, ensuring that the important information can still be fully utilized without compromising privacy. The parameter $\epsilon$, defined as the upper bound of the KL-divergence between the data distributions before and after perturbation, quantifies the effectiveness of privacy-preserving algorithms. As stated in Definition 1, a lower value of $\epsilon$ corresponds to stronger privacy guarantees. To access utility, we evaluate the F1-score of the model parameters trained from the perturbed data as a measure of the data’s effectiveness. Specifically, we train the model with the perturbed data and then validate it with the unperturbed data. The closer the validation results are to the original, the higher the utility. It is evident that as privacy protection increases (i.e. $\epsilon$ decreases), the accuracy of the trained model tends to decline. Our research seeks to strike a balance between data availability and privacy protection.

### v2

# Local Differential Privacy for Tensors in Distributed Computing Systems

@@ Abstract
Tensor-valued data, increasingly common in distributed big data applications like autonomous driving and smart healthcare, poses unique challenges for privacy protection due to its multidimensional structure and the risk of losing critical structural information. Traditional local differential privacy methods, designed for scalars and matrices, are insufficient for tensors, as they fail to preserve essential relationships among tensor elements. We introduce TLDP, a novel LDP algorithm for Tensors, which employs a randomized response mechanism to perturb tensor components while maintaining structural integrity. To strike a better balance between utility and privacy, we incorporate a weight matrix that selectively protects sensitive regions. Both theoretical analysis and empirical findings from real-world datasets show that TLDP achieves superior utility while preserving privacy, making it a robust solution for high-dimensional tensor data.
@@
Local differential privacy, tensor privacy, randomized response
@@ I Introduction
High dimensional tensor data are passionately required in a diverse array of distributed big data applications, such as autonomous driving [1] and smart healthcare [2]. In these systems, tensor data that contains sensitive private information, such as location data, facial images, or medical records, is frequently transmitted between different parties, which creates a significant risk of privacy leakage. For example, through the intercepted tensor-valued features or model parameters (in federated systems), an attacker can easily infer whether a data type was used for training the model. Typically, Local Differential Privacy (LDP) constrains an attacker’s ability to infer details about individual data from the shared data.
The conventional approach to dealing with tensors involves treating them either as a set of scalars or as a group of vectors to ensure local differential privacy. However, unlike scalar or vector data, tensor data frequently maintains the inherent structure of data, such as multimedia content and graph-based information. For instance, video data contains correlated sequential frames, while graph data consists of critical structural and relational information. Critical structural information of the tensor-valued data and interrelationships among elements could be lost by flattening them into plain vectors or treating them as scalars when protecting privacy. For example, a time-series tensor like videos in computer vision tasks or financial data in stock market analysis applications contain crucial instructional information, and flattening them goes against common sense; or consider a symmetric tensor like Diffusion Tensor Imaging (DTI) in medical imaging or the polarizability tensor in chemistry, if each element is perturbed using traditional methods, such as adding independent and identically distributed noise, the probability that the tensor still remains symmetric is extremely low. In this study, we establish a formal definition of LDP for tensors and develop approaches to satisfy this definition.
Developing schemes for tensors to achieve local differential privacy is very challenging due to the complexity of high-dimensional data. Generally, the noise introduced to the raw data is deeply affected by the sensitivity of the query function, one key factor of which is the size of the tensor. Essentially, the larger the specifications of the data tensors, the more substantial the noise tends to be, potentially leading to a decrease in the accuracy of the results. Extensive research explored LDP schemes for scalars [3, 4, 5], vectors [6, 7], and matrices [8, 9] by either adding random noise or randomly perturbing individual inputs to achieve local differential privacy and achieved promising results, but an LDP approach applicable to tensor-valued data in distributed computing systems has yet to be established. Although a few research studies like [10] introduce an LDP for tensors for distribution estimation, they cannot be directly applied for locally differential privacy protection in distributed computing systems. It is because although the perturbed tensor-valued data still retain certain statistical properties, the model analysis results on such data will be seriously affected. Therefore, in this work, we aim to tackle this challenge by developing a local differential privacy mechanism that ensures the effectiveness of the subsequent data analytics despite the high-dimensionality of tensors.
Both privacy and utility play a critical role when practitioners apply LDP in practical distributed computing systems. However, designing an LDP that balances the privacy and utility of data presents a significant challenge due to the inherent trade-off between the two. This issue is even more pronounced with tensors, where excessive noise/perturbation is usually added due to the high dimensionality of tensors, which can severely diminish their usefulness. To mitigate such degradation, various methods have been explored, as seen in [11, 12, 13]. However, most of them are for scalars or vectors instead of tensors, lacking a utility guarantee for tensors in distributed computing systems. In contrast, our approach ensures practical utility while satisfying $\epsilon-$LDP in distributed computing systems.
In this paper, we formalize the definition of LDP for tensor-valued data and innovatively introduce a Local Differential Privacy algorithm designed for Tensors of varying dimensions named TLDP, preserving the tensor-valued data’s structural information. TLDP incorporates locally differentially private and tensor-shaped noise into the tensors by randomly selecting components of the tensor with a randomized response mechanism to accomplish local differential privacy. The key idea behind it is to randomly choose whether or not to add Laplace/Gaussian noise to the tensor data so that the total amount of noise added to the tensors is lower than directly adding noise to each entity of the tensor without breaking the condition of satisfying $\epsilon-$LDP. We rigorously prove that TLDP satisfies local differential privacy both when using the Laplace/Gaussian noise and has a lower expected error than previous studies.
Additionally, we observe that different parts of a tensor may have varying sensitivities in practical applications. For example, in a road scene image, the facial or license plate number information is more critical than the background for scene understanding tasks in autonomous driving applications. The data privacy of such tasks can be further improved by incorporating a precisely tailored weight matrix that indicates the areas of the tensor that are more important than others. Therefore, we design an optimized local differential privacy mechanism based on TLDP by applying varying weights to different parts of the tensor.
@@
  • We propose an algorithm that ensures $\epsilon-$LDP for tensor-valued data named TLDP. TLDP has a lower expected error and greater utility compared to prior methods, independent of the tensor’s shape. To the best of our knowledge, it is the first work that addresses local differential privacy of tensor-valued data in machine learning-based distributed computing systems.
  • We further enhance the algorithm by incorporating a weight matrix, which selectively improves the impact of perturbation on important portions of the data, thereby strengthening privacy protection.
  • We provide rigorous theoretical proof of privacy under the TLDP mechanisms and present a series of experiments performed on various tensor-valued models and datasets. The results demonstrate that our method achieves superior utility compared to other mechanisms while maintaining the same level of privacy.
@@ II Related Work
Our research aligns closely with studies in the following areas.
II-A Primitive algorithms
II-A1 Local differential privacy by adding random noise
Primitive algorithms refer to those that protect data privacy by randomly adding random noise to individual data. They ensure privacy independently and do not rely on any external mechanisms. They include the Gaussian mechanism [14], Laplace mechanism [15], Matrix Variate Gaussian (MVG) [16], and Independent Directional Noise (IDN) [17].
Although the mechanisms mentioned above do not directly apply to tensor-valued queries, our research is nevertheless grounded in the fundamental principle of preserving privacy through the injection of noise. The Gaussian mechanism introduces independent and identically distributed (i.i.d.) Gaussian noise, calibrated according to the $l_{2}$-sensitivity of the query, guarantees $(\epsilon,\delta)$-differential privacy. Similarly, by incorporating noise derived from the Laplace distribution, scaled according to the query function’s $l_{1}$-sensitivity, the Laplace mechanism enforces strong $\epsilon$-differential privacy. The MVG mechanism is tailored for matrix-valued queries and incorporates matrix-valued noise to ensure $(\epsilon,\delta)$-differential privacy. The $l_{2}$-sensitivity of the MVG mechanism is measured using the Frobenius norm, which quantifies the discrepancy between two neighboring matrices. The IDN mechanism achieves differential privacy by appending a noise tensor of equal magnitude to the data tensor. However, LDP demands a higher level of privacy than traditional differential privacy, necessitating more noise to achieve comparable privacy guarantees, which can significantly degrade data utility due to extensive perturbation. Our mechanism is specifically tailored to address the complexities of tensor data. We demonstrate that our mechanism attains $\epsilon$-local differential privacy for tensors, as detailed in Section III. This innovation bridges the gap between traditional local differential privacy mechanisms and the sophisticated requirements of high-dimensional data, preserving privacy while maintaining utility.
Beyond these fundamental mechanisms, several sophisticated approaches have also been introduced. Mechanisms that derive their privacy guarantees from the above mechanisms are well-established, as exemplified by the composition schemes referenced in [18, 19, 20, 21, 9, 7, 6]. Among these, [18] presents an innovative accounting approach for the Gaussian mechanism that minimizes the total additive noise while preserving the same level of privacy protection. In contrast, Kairouz et al. [22] propose an optimal composition scheme that is applicable to a broader range of noise distributions. The studies in [19, 21] provide dynamic accounting methods that adjust based on the algorithm’s runtime convergence. Moreover, [20] ensures a global differential privacy guarantee in environments where datasets are continuously expanding. [9] applies differential privacy to high-order, high-dimensional sparse tensors within the IoT transmission context through individual randomized responses. However, when the central server is required to receive tensor information directly from users, this approach risks disclosing the user’s original information with a certain probability, thereby undermining the purpose of privacy protection. [7] introduced a sparse weight matrix that computes the non-zero elements of the matrix, thereby reducing the computational load of LDP when applied to large-scale datasets. Furthermore, [6] proposed a new privacy budget recovery mechanism, called ChainDP, which adds noise sequentially, so that the noise added by the previous user can be used by subsequent users, thereby reducing the overall noise level and estimation error.
II-A2 Local differential privacy by randomized response to a query
Apart from adding noise to individual data, random response is another popular approach to achieve local differential privacy. Stanley L. Warner et al. [3] introduced the randomized response mechanism in 1965, which is a technique specifically designed to preserve respondent privacy, commonly referred to as $w$-RR. The key idea is to introduce randomness into the response process, allowing respondents to answer sensitive questions without fear of exposure, while still enabling researchers to estimate the true distribution of responses. The $w$-RR technique is specifically designed for discrete datasets with exactly two distinct values. To extend its applicability, two primary paths for improvement can be explored. The first involves encoding and transforming the variable’s multiple values, as demonstrated in methods such as Randomized Aggregatable Privacy-Preserving Ordinal Response (RAPPOR) [4] and S-Hist [23], to ensure compatibility with the binary nature expected by $w$-RR. The RAPPOR employs Bloom filters and hash functions to encode a single element into a vector, then perturb the original value through both permanent and instantaneous randomized responses. It represents a method for collecting statistical data from end-user applications while preserving anonymity and ensuring robust privacy protections. However, RAPPOR has two main drawbacks: (1) the transmission overhead between the user and the data aggregator is relatively high, as each user is required to send a vector whose length is determined by the size of its bloom filter; (2) the data collector must pre-collect a list of candidate strings for frequency counting.
@@ II-A2 Local differential privacy by randomized response to a query
To address the first issue, the S-Hist method [23] offers a solution in which each user encodes the string, randomly selects a bit, applies random response techniques to perturb it, and then sends it to the data collector. This approach significantly reduces the transmission cost by minimizing the amount of data each user needs to send. For the second problem, building upon the RAPPOR-based encoding-decoding framework, Kairouz et al. [24] further proposed the $O$-RR method by introducing hash mapping and grouping operations. Hash mapping allows the method to focus on the encoded values rather than the original strings, eliminating the need to pre-collect a list of candidate strings. Additionally, the use of grouping operations reduces the probability of hash mapping value collisions. This combination of techniques enhances the robustness and flexibility of differential privacy in various data collection scenarios.
To overcome the second limitation of $w$-RR, one approach is to improve the distribution of the $w$-RR technique, making it applicable to variables with more than two values. This is achieved through mechanisms such as $k$-RR [22]. In the $k$-RR method, the original data is preserved with probability $p$ and is flipped to any other value with a probability of $\frac{1-p}{n-1}$, where $n$ is the number of possible values. This mechanism ensures that the perturbation is uniformly distributed across all possible values, thereby preserving the privacy of the data.
@@
II-B Learning With Differential Privacy
A range of studies explored the integration of differential privacy techniques into machine learning models, addressing various privacy-preserving objectives. These efforts encompass differentially private individual data [16, 28], model outputs [29], model parameters [19, 20, 21, 30, 31, 32, 33], and even objective functions [34, 35]. Our focus is especially on machine learning applications in distributed systems, which often involve frequent transmission of high-dimensional or complex tensor-valued data like raw private data, extracted features, or model parameters.
In summary, we consider our paper introduces a primitive local differential privacy scheme, which fits well in distributed computing systems but certainly can be applied in any other machine learning systems, where individual users’ tensor-valued data is private and requires to be preserved.
@@ III Preliminaries
In this section, we aim to provide our readers with foundational knowledge to facilitate the comprehension of our work.
@@ III-A Local Differential Privacy
Local differential privacy (LDP) enables data collection and analysis while ensuring strong privacy protection by perturbing data before it is shared with a central aggregator. This ensures that even the data collector cannot determine an individual’s true value with certainty. The privacy guarantee of LDP is expressed by bounding the logarithmic ratio of the output probabilities for any two possible inputs from a single user, ensuring that the private mechanism $M$ does not reveal too much about the user’s true input. This local and individualized privacy guarantee makes LDP particularly suitable for scenarios where users do not trust the data collector and need to protect their data before sharing it. Formally, given any two tensor-valued inputs from the same user $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, $y$ be the output and $M$ be the private scheme, the following condition holds:
Definition 1 ($\epsilon$-LDP).
An randomized algorithm M satisfies $\epsilon$-LDP (where $\epsilon\geq 0$), if for any tensor-valued inputs $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ from the same user, we have
$\forall y\in Range(\textbf{M}),Pr[\textbf{M}(\mathcal{X}_{1})=y]\leq e^{\epsilon}Pr[\textbf{M}(\mathcal{X}_{2})=y]$
where $Range(\textbf{M})$ represents a set of all possible outputs of the algorithm M.
Problem Definition and Notations. We suppose that in this article, each user has a number of tensor-valued data, and $\mathcal{X}\in\mathbf{R}^{I_{1}\times I_{2}\times\dots\times I_{N}}$ is any one of them. We utilize $\Delta$ to represent the range of the values of a tensor in a dataset and $[\Delta]$ to denote all possible values of the tensor. For example, $\Delta=256$ and $[\Delta]=\{0,1,2,\dots,255\}$ for an image of the MNIST dataset because that range of the values of an image of the MNIST dataset is 0-255. Without compromising generality, we assume the input domain is $[\Delta]$, i.e., $\mathcal{X}_{i_{1}i_{2}\dots i_{N}}\in[\Delta]$. If the tensor takes continuous values, it is recommended to discretize it initially. The notations used in this paper are summarized in Table I for easy reference.
@@
$I$ Product of the tensor’s dimensions, i.e., $I=I_{1}I_{2}\cdots I_{N}$
@@
$\Delta$ Range of the values of a tensor in a dataset
@@
Our goal is to perturb the original private tensor before sharing it with other parties, ensuring that users’ sensitive information can be properly preserved without compromising data utility. The parameter $\epsilon$, defined as the upper bound of the KL-divergence between the data distributions before and after perturbation, quantifies the effectiveness of privacy-preserving algorithms. As stated in Definition 1, a lower value of $\epsilon$ corresponds to stronger privacy guarantees. To access utility, we evaluate the F1-score of the model parameters trained from the perturbed data as a measure of the data’s effectiveness. Specifically, we train the model with the perturbed data and then validate it with the unperturbed data. The closer the validation results are to the original, the higher the utility. It is evident that as privacy protection increases (i.e., $\epsilon$ decreases), the accuracy of the trained model tends to decline. Our research seeks to strike a balance between data utility and privacy protection.